Another quiz question:
Find an equation of the line containing (-4, -3) and perpendicular to the line 2x-7y=14.

Answer by scott8148(6628)

perpendicular means the solpes are negative reciprocals __ so m=-7/2

using point-slope __ y-(-3)=(-7/2)(x-(-4)) __ y+3=(-7/2)x-14 __ y=(-7/2)x-17